# Exact Solution of Bipartite Fluctuations in One-Dimensional Fermions

### v1

@@ Abstract
Figure 4: Numerical results for the variance $\sigma_{M}^{\prime\prime}(t)^{2}$ with $M=40$. The number of the lattice is $1202L=$ 120. We numerically solve the Schrödinger equation using the TEBD method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019) with the maximal bond dimension $D\lesssim 2600$, calculating the variances for $U=0,-1,-2$, and $-3$. In the language of the XXZ spin-1/2 chain, our model for $U=-1,-2$, and $-3$ corresponds to a gapless regime, a critical point, and a gapped regime. The dashed line denotes the leading term of Eq. (6) for $t\gg 1$, and the dotted and dot-dashed lines are curves for $t^{2/3}$ and $t^{1/5}$, respectively. We insert these curves with the exponents $2/3$ and $1/5$ for reference.
@@
Numerical study for the variance growth of the interacting fermions.— We numerically study the interaction effect on the time evolution of the variance. Our numerical method is the time-evolving decimation method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019) with Eq. (1) and the open boundary condition ($\hat{a}_{L}=0$). We set $L$ to be $60$, and compute the variance for $U=0,-1,-2$, and $-3$. In the language of the spin-1/2 XXZ chain, $U=-2$ corresponds to the critical point at which the model is identical to the XXX chain Franchini (2017). To weaken the boundary effect, we use the variance $\sigma^{\prime\prime}_{M}(t)^{2}\coloneqq\langle(\hat{N}_{M}^{\prime})^{2}\rangle_{t}-\langle\hat{N}_{M}^{\prime}\rangle_{t}^{2}$ with $M=40$.
@@
Conclusions and prospects.— We theoretically studied the time evolution for the variance of the particle numbers in the subsystem of the fermions on the one-dimensional lattice, quantitatively explaining the recent experimental result of Ref. Wienand et al. (2023). The initial state used in this work was the alternating state. In the noninteracting fermions, we analytically derived the exact and simple expression of the variance, elucidating that the variance linearly grows in the long-time dynamics. Developing the exact analysis, we incorporated the incompleteness of the initial alternating state into our analytical expression of the variance, finding that the improved expression shows the reasonable agreement with the experimental result of Ref. Wienand et al. (2023) without any fitting parameters. Finally, our numerical results based on the TEBD method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019) find that the presence of the interaction qualitatively alters the variance growth.
As a prospect, it is intriguing to study the variance growth in other models accompanying quantum phase transitions, such as the transverse Ising model and the Su-Schrieffer-Heeger model. These models are mapped into noninteracting fermions, and one may analytically study the variance growth from the perspective of the universal nature of the phase transitions. Furthermore, studying the variance of other physical quantities, such as energy and spins, is fascinating. In another direction, it is worth studying the variance growth in open quantum systems. Depending on kinds of interactions with environments, an open quantum system approaches a nonequilibrium stationary state that is completely different from the equilibrium state addressed in this work. Thus, uncovering features of the variance growth in such a case is fundamentally interesting.
@@ Acknowledgements.
KF is grateful to Ryusuke Hamazaki and Yuki Kawaguchi for fruitful discussions and comments on the manuscript, Masaya Kunimi for helpful discussions on TEBD with a conserved quantity, and Hiroki Moriya for helpful discussions. KF and ST are grateful to Monika Aidelsburger, Immanuel Bloch, Alexander Impertro, Simon Karch, Christian Schweizer, and Julian F. Wienand for sharing the experimental data of Ref. Wienand et al. (2023) used in Fig. 3. The work of KF has been supported by JSPS KAKENHI Grant No. JP23K13029. The work of TS has been supported by JSPS KAKENHI Grants No. JP21H04432, No. JP22H01143.
@@
Department of Physics, Tokyo Institute of Technology, 2-12-1 Ookayama, Meguro-ku, Tokyo 152-8551, Japan
@@
  • (VII) Numerical perturbative calculation for the disorder potential.
@@
In the fourth line, we use the Wick theorem.

### v3

@@ Abstract
Figure 4: Numerical results for the variance $\sigma_{M}^{\prime\prime}(t)^{2}$ with $M=40$. The number of the lattice is $1202L=$ 120. We numerically solve the Schrödinger equation using the TEBD method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019), calculating the variances for $U=0,-1,-2$, and $-3$. The dashed line denotes the leading term of Eq. (6) for $t\gg 1$, and the dotted and dot-dashed lines are curves for $t^{2/3}$ and $t^{1/5}$, respectively. We insert these curves with the exponents $2/3$ and $1/5$ for reference.
Figure 3 compares our theoretical results with the experimental data of Ref. Wienand et al. (2024). Our numerical result with $\Delta=0$ can capture the variance growth in the early stage of the dynamics, but the deviation from the experimental data grows in time. On the other hand, when we turn on the random potential with $\Delta=1$, the numerical result well reproduces the experimental data. The saturation of the variance for $t\gtrsim 2$ is caused by the finite size effect and the disorder potential. These behaviors were reported in the numerical simulations of Ref. Wienand et al. (2024). We also implement a numerical perturbative calculation with $\Delta$, finding that the 6th-order perturbative result well describes the time evolution before the saturation (see Sec. VII of SM for the detailed numerical method). Comparing these results, one can see that the disorder effect is irrelevant in $t\lesssim 1$.
The dotted and dashed lines in Fig. 3 show the analytical results for the linear growth of Eqs. (6) and (11). We find that the latter including the incompleteness of the initial alternating state exhibits the good agreement with the experimental linear growth, while the former not including the incompleteness systematically deviates from the experimental data. We here emphasize that our analytical result (11) agrees well with the experimental data without any fitting parameters under the assumption of (7). Note that one can see that the disorder effect emerges before the finite $M$ effect by comparing the numerical data for $\Delta=0$ and 1 with Eq. (11), though the localization length is larger than $M=8$ Wienand et al. (2024).
We comment on the fact that our incomplete initial density matrix of Eq. (7) cannot describe superpositions of quantum states between different sites. The experiment of Ref.  Wienand et al. (2024) is expected to suppress such entangled states since deep optical lattices are used for the initial state preparation, but we numerically study the effects of the superpositions on the variance growth by extending Eq. (7) to a more general state. As illustrated in Sec. VIII of SM SM , we numerically confirm that the results of Eq. (7) show better agreement with the experiment than ones of the general state. This finding strengthens the justification for using Eq. (7).
Finally, we discuss a time scale for local equilibration on the basis of our exact solution of Eq. (9) for a finite $M\gg 1$, where the linear growth of the subtracted variance, namely $2\left(n_{\rm even}-n_{\rm odd}\right)^{2}t/\pi$, becomes dominant (see Eq. (11)). When taking the long-time limit of Eq. (9), we obtain the stationary value of the subtracted variance $\lim_{t\rightarrow\infty}(\sigma_{M}(t)^{2}-\delta\sigma_{M}(t)^{2})=(n_{\rm even}-n_{\rm odd})^{2}M/4$. Then, we can estimate the local equilibration time $T_{*}$ by equating the expressions for these two subtracted variances, finding $T_{*}\sim\pi M/8$. Interestingly, our exact solution leads to the fact that $T_{*}$ is independent of the initial value of $n_{\rm even}-n_{\rm odd}$.
@@
Numerical study for the variance growth of the interacting fermions.— We numerically study the interaction effect on the time evolution of the variance. Our numerical method is the TEBD method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019) with Eq. (1) and the open boundary condition ($\hat{a}_{L}=0$). We set $L$ to be $60$, and compute the variance for $U=0,-1,-2$, and $-3$. In the language of the XXZ chain, $U=-2$ corresponds to the critical point at which the model is identical to the XXX chain Franchini (2017). To weaken the boundary effect, we use the variance $\sigma^{\prime\prime}_{M}(t)^{2}\coloneqq\langle(\hat{N}_{M}^{\prime})^{2}\rangle_{t}-\langle\hat{N}_{M}^{\prime}\rangle_{t}^{2}$ with $M=40$. The dependence on the bond dimensions and the time resolutions of our numerical calculations is described in Sec. X of SM SM .
@@
Finally, we study the effect of the random potential on the variance growth. We numerically find that the disorder effect emerges in the early stage of the dynamics and the power-law-like growth in Fig. 4 becomes unclear, as described in Sec. IX of SM SM . This numerical finding suggests that one needs to decrease the strength of the random potential to observe the power-law-like growth in the interacting fermions, namely the XXZ model.
Future prospects.— Our theoretical results, based upon the exact solutions of the noninteracting fermions and the numerical findings of the interacting fermions will stimulate further research on emergent hydrodynamics and local equilibration of quantum many-body systems. As a prospect, it is fascinating to analytically comprehend the interaction effect via generalized hydrodynamics Castro-Alvaredo et al. (2016); Bertini et al. (2016); Bulchandani et al. (2017, 2018); Doyon and Yoshimura (2017); Doyon et al. (2017, 2018); Collura et al. (2018); De Nardis et al. (2018); Schemmer et al. (2019); Gopalakrishnan and Vasseur (2019); Doyon (2020); Alba et al. (2021); Malvania et al. (2021); Bouchoule and Dubail (2022); Essler (2023) and ballistic macroscopic fluctuation theory Doyon et al. (2023a, b). In another direction, it is worth studying the variance growth in open quantum systems. Depending on the kinds of interactions with environments, an open quantum system approaches a nonequilibrium stationary state being different from the equilibrium state addressed in this work. Thus, uncovering features of the variance growth in such a case is fundamentally intriguing.
@@ Acknowledgements.
KF is grateful to Ryusuke Hamazaki and Yuki Kawaguchi for fruitful discussions and comments on the manuscript, Masaya Kunimi for helpful discussions on TEBD with a conserved quantity, and Hiroki Moriya for helpful discussions. KF and ST are grateful to Monika Aidelsburger, Immanuel Bloch, Alexander Impertro, Simon Karch, Christian Schweizer, and Julian F. Wienand for sharing the experimental data of Ref. Wienand et al. (2024) used in Fig. 3 of the main text and Figs. S-3 and S-4 of SM SM . The work of KF has been supported by JSPS KAKENHI Grant No. JP23K13029. The work of TS has been supported by JSPS KAKENHI Grants No. JP21H04432, No. JP22H01143.
Note added.— After uploading our manuscript on arXiv, O. Gamayun have reported that the bipartite fluctuations growing from the alternating state is related to that from the domain-wall state Gamayun (2024).
@@
Department of Physics, Institute of Science Tokyo, 2-12-1 Ookayama, Meguro-ku, Tokyo 152-8551, Japan
@@
  • (VII) Numerical perturbative calculation for the disorder potential,
  • (VIII) Effect of the superposition of the initial state in the noninteracting fermions,
  • (IX) Effect of the random potential on the interacting fermions,
  • (X) Dependence on the bond dimension and the time resolution in the numerical calculations.
@@
In the fourth line, we use the Wick theorem. This kind of determinantal form was derived in several previous works Schönhammer (2007); Eisler and Rácz (2013); Parez et al. (2021).
@@ VII Numerical perturbative calculation for the disorder potential
VIII Effect of the superposition of the initial state in the noninteracting fermions
We numerically study the variance growth in the noninteracting fermions using a generalized incomplete initial state, which can include the superpositions between different sites. The setup is the same as the one used in Fig. 3 of the main text except for the initial state. We here use the generalized initial density matrix as an initial state, which is defined by
$\hat{\rho}_{\rm two{\text{-}}site}$ $\coloneqq\dfrac{1}{\Xi}\exp\left(-\beta\hat{H}_{\rm two{\text{-}}site}+\beta\mu\hat{N}_{\rm tot}\right).$ (S-81)
Here, the Hamiltonian $\hat{H}_{\rm two{\text{-}}site}$ and the total number operator $\hat{N}_{\rm tot}$ are given by
$\hat{H}_{\rm two{\text{-}}site}$ $\coloneqq-J\sum_{j=-L/2}^{L/2-1}\left(\hat{a}^{\dagger}_{2j+1}\hat{a}_{2j}+\hat{a}^{\dagger}_{2j}\hat{a}_{2j+1}\right)+\sum_{j=-L/2}^{L/2-1}\left(V_{\rm odd}\hat{a}^{\dagger}_{2j+1}\hat{a}_{2j+1}+V_{\rm even}\hat{a}^{\dagger}_{2j}\hat{a}_{2j}\right),$ (S-82)
$\hat{N}_{\rm tot}$ $=\sum_{j=-L}^{L-1}\hat{a}^{\dagger}_{j}\hat{a}_{j}$ (S-83)
with the two-site hopping parameter $J$, the even-site potential $V_{\rm even}$, and the odd-site potential $V_{\rm odd}$. In the following, we set $\beta=1$ without loss of generality. For a given $J$, the values of $V_{\rm even}$, $V_{\rm odd}$, and $\mu$ are numerically determined such that $n_{\rm even}$ and $n_{\rm odd}$ satisfy the following conditions:
$I=\dfrac{n_{\rm odd}-n_{\rm even}}{n_{\rm odd}+n_{\rm even}}=0.91,~{}~{}~{}~{}\nu=\dfrac{n_{\rm odd}+n_{\rm even}}{2}=0.44.$ (S-84)
For $J=0$, this generalized initial state is identical to the initial density matrix used in the main text. Under this situation, we can study the effect of the initial superposition on the variance growth in the noninteracting fermions.
Figure S-3: Numerical results for the dependence of the initial variance with the generalized initial state of Eq. (S-81) on the subsystem size $M$ and the initial hopping parameter $J$. The circle, pentagon, and square markers represent the initial variance obtained numerically with $J=0,1,$ and $2$, respectively. The red hexagon markers with the bar is the experimental data for Fig. S8 of Ref. Wienand et al. (2024). For the comparison with the experiment, we plot the variance $2\sigma^{\prime}_{M}(t)^{2}$ for the numerical results and $\sigma_{\rm exp}(t)^{2}$ for the experimental result because the experiment studies two-ladder lattice systems and thus the observed variance is twice as large as ours. This graph clearly shows that the deviations between the experimental and numerical results become large when the initial hopping parameter $J$ increases.
Figure S-3 displays the dependence of the initial variance on the subsystem size $M$ and the hopping parameter $J$, clearly showing that the experimental date is very close to the numerical data with $J=0$. We also plot the time evolution of the variance growth with the generalized initial state of Eq. (S-81) in Fig. S-4. Then, we find the numerical results with $J=0$ exhibit better agreement with the experimental data of Ref. Wienand et al. (2024). All these numerical findings suggest that the initial superpositions of the experiment is strongly suppressed, as pointed out in the experimental paper of Ref. Wienand et al. (2024).
Figure S-4: Numerical results for the variance growth in the noninteracting fermions with the generalized initial state of Eq. (S-81). The figures display the time evolution of the variance growth for (a) $M=14$, (b) $M=8$, and (c) $M=2$. For the comparison with the experiment, we plot the variance $2\sigma^{\prime}_{M}(t)^{2}$ for the numerical results and $\sigma_{\rm exp}(t)^{2}$ for the experimental results because the experiment studies two-ladder lattice systems and thus the observed variance is twice as large as ours. The numerical results with $J=0$ exhibit better agreement with the experimental data for Fig. S8 of Ref. Wienand et al. (2024) compared to the numerical data with the finite hopping parameter $J\neq 0$.
IX Effect of the random potential on the interacting fermions
We numerically study how the random potentials affect the variance growth in the interacting fermions. The setup in this section is the same as the one used in Fig. 4 of the main text, but we use the Hamiltonian $\hat{H}_{\rm random}$ with the random potential, which is defined by
$\hat{H}_{\rm random}\coloneqq-\sum_{j=-L}^{L-1}\left(\hat{a}_{j+1}^{\dagger}\hat{a}_{j}+\hat{a}_{j}^{\dagger}\hat{a}_{j+1}\right)+\sum_{j=-L}^{L-1}V_{j}\hat{n}_{j}+U\sum_{j=-L}^{L-1}\hat{n}_{j+1}\hat{n}_{j},$ (S-85)
where the random potential $V_{j}$ is independently sampled from the uniform distribution with the range $[-\Delta,\Delta]$. Here, $\Delta$ is the real positive parameter. We solve the Schrödinger equation with Eq. (S-85) via the TEBD method, numerically investigating the effect of the random potential on the variance growth. In what follows, the time evolution of the variance is computed by using 200 samples for the ensemble average of the random potential.
Figure S-5: Numerical results of the variance growth in the interacting fermions with the random potential for (a) $U=-1$, (b) $U=-2$, and (c) $U=-3$. In each graph, we put the data with the strength $\Delta=0$ and $1$ of the random potential to make visible the disorder effect clearly. The numerical data with $\Delta=0$ is the same as the data of Fig. 4 of the main text.
Figure S-5 shows the numerical results for the variance growth with $\Delta=0$ and $1$. Our numerical simulations for the interaction parameter $U=-1$ and $-2$ shows that the disorder effect becomes large as time goes by. Thus, in these cases we expect that the power-law-like growth seen in the interacting fermions with $\Delta=0$ will be strongly disturbed in the long-time dynamics. On the other hand, the disorder effect for $U=-3$ is not large in the time region we can numerically access.
X Dependence on the bond dimension and the time resolution in the numerical calculations
Figure S-6: Dependence on the bond dimension $D$ and the time resolution $dt$ for the variance growth in the interacting fermions with (a) $U=0$, (b) $U=-1$, (c) $U=-2$, and (d) $U=-3$.
We describe the details of our numerical simulation based on the TEBD method. The notations used in this section are the same as those used in Fig. 4 of the main text. In the TEBD method, we usually express a quantum state |$\ket{\psi}$ as
|$\ket{\psi}=\sum_{\alpha_{1}=1}^{D_{1}}\sum_{n_{1}=0}^{1}\cdots\sum_{\alpha_{1}=1}^{D_{L-1}}\sum_{n_{L-1}=0}^{1}\sum_{n_{L}=0}^{1}\Gamma_{\alpha_{1}}^{[1],n_{1}}\lambda_{\alpha_{1}}^{[1]}\Gamma_{\alpha_{1},\alpha_{2}}^{[2],n_{2}}\lambda_{\alpha_{2}}^{[2]}\cdots\Gamma_{\alpha_{L-2},\alpha_{L-1}}^{[L-1],n_{L-1}}\lambda_{\alpha_{L-1}}^{[L-1]}\Gamma_{\alpha_{L-1}}^{[L],n_{L}}$ (S-86)
with the real variable $\lambda$ and the complex variable $\Gamma$, where their subscripts are omitted for brevity. The parameter $D_{j}~{}({j\in\{1,...,L-1\}})$ is referred to as a bond dimension and plays the role of a cutoff parameter in the numerical simulation. Here, we emphasize that our model conserves the total particle number but the expression of Eq. (S-86) does not incorporate the merit of such conservation law. In order to utilize the conservation law, it is helpful to consider the following expression instead of Eq. (S-86):
|$\ket{\psi}=\sum_{q_{1}}\sum_{\alpha_{1}=1}^{D_{1}(q_{1})}\sum_{n_{1}=0}^{1}\cdots\sum_{q_{L-1}}\sum_{\alpha_{1}=1}^{D_{L-1}(q_{L-1})}\sum_{n_{L-1}=0}^{1}\sum_{q_{L}}\sum_{n_{L}=0}^{1}\Gamma_{\alpha_{1}}^{[1],n_{1},q_{1}}\lambda_{\alpha_{1}}^{[1],q_{1}}\Gamma_{\alpha_{1},\alpha_{2}}^{[2],n_{2},q_{2}}\lambda_{\alpha_{2}}^{[2],q_{2}}\cdots\Gamma_{\alpha_{L-2},\alpha_{L-1}}^{[L-1],n_{L-1},q_{L-1}}\lambda_{\alpha_{L-1}}^{[L-1],q_{L-1}}\Gamma_{\alpha_{L-1}}^{[L],n_{L},q_{L}},$ (S-87)
where we introduce the new integer $q_{j}~{}(j\in\{1,...,L\})$ to reduce the computational costs. The critical point of this expression is that the bond dimension $D_{j}(q_{j})$ depends on $q_{j}$. In our numerical calculation, we introduce the maximal bond dimension $D$ such that all the bond dimensions $D_{j}(q_{j})$ are smaller than $D$. Note that we need a larger bond dimension when we do not consider the particle-number conservation in the numerical calculation. As to the time discretization, we use the 2nd-order checkerboard discretization based on the Suzuki-Trotter decomposition with the time resolution $dt$.
Figure S-6 displays the time evolution of the variance corresponding to Fig. 4 of the main text. Then, one can see that the numerical data show almost the same behavior irrespective of the values of $D$ and $dt$. Hence, our numerical results are well convergent. In Fig. 4 of the main text, we use the numerical data with the values of $D$ and $dt$ summarized in Table S-1.
Table S-1: Values of $D$ and $dt$ used for the numerical data of Fig. 4 of the main text
Interaction parameter $U$ Time resolution $dt$ Bond dimension $D$
0 0.05 350
-1 0.05 350
-2 0.05 400
-3 0.05 400